a boat takes 3 hours to go 24 miles upstream.it can go 36 miles downstream in the same time,Find the speed of the current and the speed of the boat in still water if x=the speed of the boat in still water and y=the speed of the current.
Answer by mananth(16946)

a boat takes 3 hours to go 24 miles upstream.it can go 36 miles downstream in the same time,Find the speed of the current and the speed of the boat in still water if
x=the speed of the boat in still water
y=the speed of the current.
downstream speed = x+y
upstream speed = x-y
...
time downstream = 36/(x+y)=3
3x+3y=36
time upstream = 24/(x-y)=3
3x-3y=24
add the equations
6x=60
/6
x= 10 mph speed of boat
...
plug value of x in equation
30+3y=36
3y=36-30
3y=6
y=2 mph speed of current
...
m.ananth@hotmail.ca